OpenStudy (anonymous):

Can someone help me with this please?? 3x^2 +1x-2/3x^2+10x-8

OpenStudy (aum):

What is the question?

OpenStudy (anonymous):

would it be 6+1x-2 and then 6+10x-8??

OpenStudy (aum):

You are still not telling us what you are asked to do. Are you asked to factor it?

OpenStudy (mathmath333):

\(\large\tt \color{black}{\dfrac{3x^2 +1x-2}{3x^2+10x-8}}\) \(\large\tt \color{black}{=\dfrac{3x^2 +3x-2x-2}{3x^2+12x-2x-8}}\) \(\large\tt \color{black}{=\dfrac{3x(x +1)-2(x+1)}{3x(x+4)-2(x+4)}}\) \(\large\tt \color{black}{=\dfrac{(3x-2)(x+1)}{(3x-2)(x+4)}}\) \(\large\tt \color{black}{=\dfrac{(x+1)}{(x+4)}}\)

OpenStudy (anonymous):

3x^2+ 1x-2 / 3x^2+10x-8

OpenStudy (anonymous):

just simplify

OpenStudy (anonymous):

@mathmath333 can you tell me how you did the work please?

OpenStudy (mathmath333):

i factored numerator and denominator , do u know how to factor quadratic equations

OpenStudy (anonymous):

no

OpenStudy (mathmath333):

this one is also good http://www.mathsisfun.com/algebra/factoring-quadratics.html

OpenStudy (mathmath333):

\(\large\tt \color{blue}{\text{there is also another method }}\) \(\large\tt \color{black}{ax^2+bx+c=(x-\alpha)(x-\beta)}\) \(\large\tt \color{black}{where~~\alpha=\dfrac{-b+\sqrt{b^2-4ac}}{2a}}\) \(\large\tt \color{black}{where~~\beta=\dfrac{-b-\sqrt{b^2-4ac}}{2a}}\)

OpenStudy (anonymous):

Can it be simplify more @mathmath333

OpenStudy (mathmath333):

no my answer is simplified atmost